OpenStudy (fabgirl1994):

If I get this wrong, I fail ?! Algebra 1 question? I have the first part.. just need the second! Help!

OpenStudy (fabgirl1994):

Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer.

OpenStudy (fabgirl1994):

Both Ray and Kelsay are correct because a third degree polynomial can have at most three x intercepts and always has  one y intercept. So it can have 4 intercepts and 3 zeroes. It cross the x three time and the y once or depending where the exponents are either on x or y it can cross the y three times and the x once.

OpenStudy (fabgirl1994):

Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. g(x) = x^3 − x^2 − 4x + 4 g(x) = x^3 + 2x^2 − 9x − 18 g(x) = x^3 − 3x^2 − 4x + 12 g(x) = x^3 + 2x^2 − 25x − 50 g(x) = 2x^3 + 14x^2 − 2x − 14 Create a graph of the polynomial function you selected from Question 2.

OpenStudy (fabgirl1994):

g(x) = x^3 − x^2 − 4^x + 4 g(x) is large and positive when x is large and positive; g(x) is large and negative when x is large and negative. The y-intercepts are  4,  The zeroes are -2, 2, 1

OpenStudy (fabgirl1994):

Then the last part, I made a graph but I don't need to post it... so that information above will help with this next part that I need help on.

OpenStudy (fabgirl1994):

Part B The second part of the new coaster is a parabola. Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros. The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan. Create a graph of the polynomial function you created in Question 4.

OpenStudy (shaik0124):

@zaynab123 check out this question

OpenStudy (fabgirl1994):

Everything before PART B is done! There's 3 questions (the last is the graph that I didn't upload) The question, then the next is the answer. That information is needed for part b. Thats what I need help on is part B.

OpenStudy (anonymous):

what numbers did you pick for a, b, and c?

OpenStudy (fabgirl1994):

A B and C see what?

OpenStudy (anonymous):

to create the parabola you need to pick numbers for a, b, and c function f(x).

OpenStudy (fabgirl1994):

You have to create a parabola in that pattern, thats what i need help on.

OpenStudy (anonymous):

They gave you the format for the function. Creating the parabola just means pick whatever numbers you want for a, b, and c. \[f(x) = ax^2+bx+c\] Also, when you're copying and pasting you questions, go in and use the ^ sign so it's clear that you're working with exponents.

OpenStudy (fabgirl1994):

I did use the ^ ? And okay, ill pick 5 9 and 2. Can you work with that? I have a time limit and it's getting low on the assignment.

OpenStudy (anonymous):

\[f(x)=5x^2+9x+2\] ok, now use the quadratic formula to find the zeros. Plug in the numbers you picked \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (fabgirl1994):

I really don't know how to do all that.

OpenStudy (anonymous):

a = 5 b = 9 c = 2 Replace the letters with the numbers